OpenStudy (anonymous):

A ball is dropped from a treetop 256 feet above the ground. How long does it take to hit the ground? (s = 16t^2) 4 seconds 8 seconds 16 seconds

OpenStudy (anonymous):

what is (s=16t^2) ??????

OpenStudy (anonymous):

\[s=16t^{2} \]

OpenStudy (anonymous):

s = distance right? lol

OpenStudy (anonymous):

Ok, I understood what they hv done, divided value of g(in ft) by 2.... In this case distance "s" is given by the formula s = ut + 1/2gt^2 here u = initial velocity = 0 (so ut = 0) g = acceleration due to gravity = 32 ft/s^2 (second^2) t = ? we have to find out So let time = t s ie. t second putting these values in the formula s = 0*t + 0.5*32*t^2 ft/s^2 256 = 16 t^2 256/16 = t^2 16 = t^2 root(16) = t 4 = t

OpenStudy (anonymous):

so ans is 4 second

OpenStudy (anonymous):

\[256=16t^2\]\[t^2=\frac{256}{16}\]\[t=\sqrt{16}=4 \]

OpenStudy (anonymous):

hey willow, pls click GOOD ANSWER button if u r satisfied with the answer :))